# Post-detection inference for sequential changepoint localization

### v4

@@ Appendix A Omitted proofs
where the last inequality follows from Assumption 3.5. Similarly,
@@ A.1 Auxilary Lemmas
Suppose $\tau_{N+1}$ is independent and identically distributed (i.i.d.) with $\tau_{1},\cdots,\tau_{N}$. By Theorem 1.A.2 of [26], $\tau_{N+1}$ is stochastically larger than $\tau$ implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\tau\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\tau_{N+1}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$. By direct calculation,
@@ Appendix C Computationally and statistically efficient subroutines in Algorithm 2
Let $L_{t}(\theta_{1};X_{1},\cdots,X_{n})$ be the likelihood of data $X_{1},\cdots,X_{n}$ under $\mathcal{H}_{0,t}$. Then, for $n\geq t$,
$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$. The point estimate of $T$ can be naturally extended to:
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (S2)
The prototypical test statistic that we employ in our simulations is the likelihood-ratio-based test statistic $M_{t}^{(n)}$, where for $t,n\in\mathbb{N},t\leq n$,
$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
@@
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists, and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}_{t}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$. In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
Let us now fix $t\in\{1,\cdots,\tau\}$. Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set. However, it is often straightforward to compute $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$. Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient. For example, we can bound it by $\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. Again, we bound $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)$ by $V_{t,L}^{j}$, where for $L<\infty$,
$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (S6)
And for $L=\infty$,
[…]
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section C.1). We define the confidence set as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (S8)
We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}$ is the set produced by Algorithm 2, since
$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
@@
Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 4.3.
Corollary C.1.
$\mathcal{C}^{\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.3.
Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section H.4.
@@ C.1 A concrete example: Gaussian mean shift problem
Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence [12] for $\theta_{1}$
@@
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (S4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.2) and they are finite. So, we can have $L=\infty$.
If $t_{2}^{j}<t$, we have $V_{t,\infty}^{j}=0$. For $t_{2}^{j}\geq t$, it can be simplified in our setting as below:
@@
$V_{t,\infty}^{j}$ $=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\max_{1\leq i\leq t^{\prime}}\frac{L_{i}\left(\frac{\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta)}{t^{\prime}-i+1};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}{L_{t}\left(\frac{\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j}}{t^{\prime}-t+1}+\theta;Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})}$
$=\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\frac{\max_{1\leq i\leq t^{\prime}}L_{i}(\bar{Y}_{i:t^{\prime}}^{j}(\theta_{t^{\prime}}^{i});Y_{1}^{t}(X_{1}^{j},\theta_{t^{\prime}}^{i}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta_{t^{\prime}}^{i}))}{L_{t}(\bar{\epsilon}_{t:t^{\prime}}^{j};\epsilon_{1}^{j},\cdots,\epsilon_{t^{\prime}}^{j})},$
where $\bar{\epsilon}_{i:t^{\prime}}^{j}=\frac{1}{t^{\prime}-t+1}\sum_{k=t}^{t^{\prime}}\epsilon_{k}^{j},\bar{Y}_{i:t^{\prime}}^{j}(\theta)=\frac{1}{t^{\prime}-i+1}\sum_{k=i}^{t^{\prime}}Y_{k}^{t}(X_{k}^{j},\theta),$ and $\theta_{t^{\prime}}^{i}$ is the unique element that maximizes $L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)$, or equivalently minimizes
$-2\log L_{i}\left(\bar{Y}_{i:t^{\prime}}^{j}(\theta);Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta)\right)=$
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta-\theta_{0})\right)^{2}+\displaystyle\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{(t-i)(\theta-\theta_{0})}{t^{\prime}-i+1}\right)^{2}+\text{const},\text{ if }i<t,\\ \text{const},\text{ if }i=t,\\ \displaystyle\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}+\theta-\theta_{0}\right)^{2}+\text{ const},\text{ if }i>t,\end{cases}$
over $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant refers to the terms that do not involve $\theta$. Note that $\theta_{t^{\prime}}^{i}$ has a simple closed-form expression since the above equation is a quadratic function of $\theta$. And hence, $V_{t,\infty}^{j}$ can be computed easily.
Remark C.2.
@@
For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, these are defined as
@@
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (S11)
@@
 $\text{ and }\penalty 10000\ \tau_{\text{wsr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}.$ (S12)
@@
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section H.3.
@@
    The coupling function $G$ is available (which is always the case when distributions are continuous and in the case of discrete distributions, even if a suitable $G$ is not directly available, our method may still be implementable, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution).
@@
    The quantities $t_{j}^{1}$ and $t_{j}^{2}$ are computable (either exactly or numerically).
  4. 4.
    If we employ the likelihood-ratio-based test statistic (S4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).

### v6

@@ Appendix A Omitted proofs
where the last inequality follows from Assumption 1. Similarly,
@@ A.1 Auxilary Lemmas
Suppose $\tau_{N+1}$ is independent and identically distributed (i.i.d.) with $\tau_{1},\cdots,\tau_{N}$. By Theorem 1.A.2 of [shaked2007stochastic], $\tau_{N+1}$ is stochastically larger than $\tau$ implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\tau\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\tau_{N+1}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$. By direct calculation,
@@ Appendix C Computationally and statistically efficient subroutines in Algorithm 2
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists, and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}_{t}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
Let us now fix $t\in\{1,\cdots,\tau\}$ and consider the test statistic $M_{t}$ of the form in (9), with some $R^{(t)}_{n}=R^{(t)}_{n}(\{X_{n}\}_{n})$ and $S^{(t)}_{n}=S^{(t)}_{n}(\{X_{n}\}_{n})$. Computing the detection time $\tau_{j,t}^{\theta}$ for each $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set. However, it is often straightforward to compute $t_{1}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j,t}$. Also, we can often derive a suitable upper bound on the threshold, which is computationally and statistically efficient. We consider the following upper bound:
$V_{t,L}^{j}:=\begin{cases}\displaystyle\max\left\{\max_{t\leq n\leq t_{1}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}R^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}\right),\max_{1\leq n\leq t-1}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}S^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}\right)\right\},\\ \hskip 341.43306pt\text{if }t\leq t_{1}^{j}\leq L;\\ -\infty,\text{ if }t>t_{1}^{j};\\ \infty,\text{ if }t\leq t_{1}^{j}\text{ and }t_{1}^{j}>L.\end{cases}$ (S13)
We consider the following confidence set:
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B}\}\text{ and}$ (S14)
We observe that $\mathcal{C}_{1}^{\prime}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the set produced by Algorithm 2, since
$\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B}\}$ $\geq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
@@
Thus, we obtain the following coverage guarantee as an immediate corollary of Theorem 7.
Corollary 1.
$\mathcal{C}^{\prime}_{1}$ defined in (S14) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 7.
Next, we present a concrete example of the Gaussian mean shift problem to clarify our method.
@@ C.1 A concrete example: Gaussian mean shift problem
Let the pre-change distribution $F_{0}$ be $N(\theta_{0},1)$ and the post-change model be $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, with $\Theta_{1}=\{\theta:\theta\geq\theta_{1}\}$ such that $\theta_{1}>\theta_{0}$.
In this setting, we have the following example of a $1-\beta$ confidence sequence [howard2021time] for $\theta_{1}$
@@
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the test statistic $M_{t}$, as defined in (9), with $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}}(X_{i})},$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w\geq 0$ being some weight function such that $\int_{\theta\in\Theta_{1}}w(\theta)d\theta=1$. Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark 5) and they are finite. So, we can have $L=\infty$. Note that, here we have
$\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}R^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}\right)=\exp\left((\theta_{0}-\theta_{1})\sum_{i=t}^{n}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{n-t+1}{2}(\theta_{1}^{2}-\theta_{0}^{2})\right),$
and
@@
$\sup_{\theta^{\prime}\in\mathcal{S}^{\prime}_{\tau-t+1}}S^{(t)}_{n}\left(\{Y_{n}^{t}(X_{n}^{j},\theta^{\prime})\}_{n}\right)=\displaystyle\underset{\theta\in\Theta_{1}}{\int}\exp\left((\theta-\theta_{0})\sum_{i=n}^{t-1}(\epsilon_{n}^{j}+\theta^{\prime}_{t,\min})+\frac{t-n}{2}(\theta_{0}^{2}-\theta^{2})\right)w(\theta)d\theta,$
where $\theta^{\prime}_{t,\min}=\inf\mathcal{S}^{\prime}_{\tau-t+1}$. And hence, $V_{t,\infty}^{j}$ in (S13) can be computed easily.
Remark 5.
@@
For weighted CUSUM or SR detectors, finding $t_{1}^{j}$ is often straightforward. For some weight $w(\theta)$ such that $\int_{\theta\in\Theta_{1}}w(\theta)=1$, these are defined as
@@
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)\geq A\right\}$ (S17)
@@
 $\text{ and }\penalty 10000\ \tau_{\text{wsr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)\geq A\right\}.$ (S18)
@@
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta_{1}\}$, for some $\theta_{1}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta_{1}\}$, for some $\theta_{1}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors.
@@
    The coupling function $G$ is available (which is always the case when distributions are continuous).
@@
    The quantities $t_{j}^{1}$ are computable (either exactly or numerically).